Prev Next

SW&IT Meeting Minutes 2020-06-08

Attendees: Carl, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20200601
LR20200601 - Bec, Matthieu 40.0
LR20200601 - Holmberg, Carl 40.0
LR20200601 - Edmeades, David 72.0
LR20200601 - Sola, Igor 80.0
LR20200601 - Hooper, Stephen 80.0
LR20200601 - Zhang, Yang 40.0
LR20200601 - Zhang, Xianyu 40.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • s/w weekly back to 10:30am

Team reports

MB

  • Last week
    • OC, Management
    • TMS/AO-cravat meetings
  • This week
    • SHARK-VIS must finish RFA
    • Applicant / Students interviews
    • DMS
      • websockets, AdSec, etc.
      • LBC stiched png preview
      • max_bytes for ovms
      • merge IMS into DMS
    • cobbler / kickstart, ims f30
    • RPM/CI short tutorial

PK

  • added following to ~/.gitconfig:
[alias]
        resign = rebase --exec 'git commit --amend --no-edit -n -S' -i
        ll = log --graph --abbrev-commit --decorate --format=format:'%C(red)%G? %C(bold blue)%h%C(reset) - %C(bold cyan)%aD%C(reset) %C(bold green)(%ar)%C(reset)%C(bold yellow)%d%C(reset)%n'' %C(white)%s%C(reset) %C(dim white)- %an %C(reset)' --all

DE

  • Macbook work
    • Data backup
    • Found out about Apple's dumb proprietary M.2 pinout
    • Ordered adapter
    • Removed swollen battery
    • After an intense battle with iFixit cart software, we got the new battery ordered
  • Saftey training
  • Work with LeRoy on summit server room sensors, looking like data are being taken
Next:
  • MB request cosmetic edit on Google Zoom rooms
  • Cisco training
  • Macbook battery swap
  • Drop off old battery for disposal
  • Chase CDW re: Veeam quote
  • Write Luci server spec comparison doc

YZ

  • Previous:
    • TMS software meeting (regular every other Thursday afternoon).
    • AO-Cravat meeting (regular every other Friday morning).
    • GCS/GCSGUI troubleshooting/patch release.
  • Next:
    • GCS/GCSGUI new patch release (for IT 8112).
    • TMS and AO-Cravat wiki cleanup and reorganization (add meeting minutes, etc.)
    • TMS and AO-Cravat software work.
    • Pickup some previous work (Olga's MODS UI, Python stuffs).
    • Safety training.

XZ

  • Last week
    • Shell rest procedure
    • Make oaa_lib backwards compatible for new4D
    • Troubleshooting of SX ASM issue
    • Write a document for IT8090
  • This week
    • AO operation support?
    • new 4D support?
    • Power Backplane action items

CH

  • Last Week:
    • AGW:
      • emailed binary encoder handing query to Omron, support engineer returns to work today.
      • JM has ordered new encoder and interface board, expected by end of June and mid-August.
    • AHS: restarted work on LBTI alarm patch.
    • ARGOS: ran build process to review steps needed to create RPM spec.
    • DMS: committed new sensor handling code to branch.
  • This Week:
    • AHS: continue work on LBTI alarm patch.
    • Argos: continue work on RPM build.
    • MODS: check if there's a more elegant way to check MODS AGW and guide camera connectivity after restarts.
    • TCS, etc: update leap second file, per IT7319.
    • Safety training

IS

Last week:
  • Safety training
  • LUCI
    • Fix on #IT7952, waiting for it to be tested.
    • LUCI WebIO device http API (ongoing)
  • PIT
    • double submission issue reported
    • Some PIT admin tool backend work
This week:
  • Safety Training
  • LUCI
    • Partner with Barry/Dave on testing #IT7952
    • work on configuration of LUCI WebIO, the HTTP GET api seems to be the same but on this device seems to be not consistent. Have to work a bit more on it. David E. provided support for network installation and backup configuration if needed.
    • Need to partner with Jenny and revert/test to old GEIRS in LUCI2 osa-luci.sim
  • PIT
    • Work on the PIT admin tool backend.
  • OT
    • Partner with Andrew to work on MODS XSLT

SH

  • Moves for observers in Tucson
    • 4 heads from 507 to 530
    • computer issue with 507-2
    • new computer for 530 and the 4 heads
  • VPN for Tucson (not much work)
  • Repository permissions (nothing heard back)
  • Spare monitors Jim and Jogn.
  • Anydesk woes.
  • iLocater requests to mountain and I
  • Server room sensors
  • TCS logs

next:

  • VPN Tucson
  • Satellite images
  • Account consolidation
Topic revision: r13 - 10 Jun 2020, CarlHolmberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback